OpenStudy (anonymous):

3=(1.02)^1t . I have the answer I just need to know the steps needed to solve the problem. I don't understand how to bring the t down.

OpenStudy (precal):

take the log of both sides or natural log of both sides.

OpenStudy (anonymous):

\[3=(1.02)^t\] \[\ln(3)=t\times \ln(1.02)\] \[t=\frac{\ln(3)}{\ln(1.02)}\]

OpenStudy (anonymous):

if you need another step it is \[\ln(3)=\ln((1.02)^t)=t\ln(1.02)\]

OpenStudy (anonymous):

but that step ,while being correct, should be unnecessary because \[b^x=e^{x\ln(b)}\] so \[(1.02)^t=e^{t\ln(1.02)}\] hence \[\ln((1.02)^t)=t\ln(1.02)\]